eviant:

PLS HELP!!! Find the slope of a line perpendicular to 2x-y=16

eviant:

@Vocaloid

Vocaloid:

first you will arrange 2x - y = 16 in y = mx + b form to find the slope then you will take the reciprocal and multiply by -1 to get the solution

eviant:

-y=2x+16

Vocaloid:

have to be careful, re-arrange doesn't just mean move the terms around, you have to perform the operations necessary to move the terms for example, for 2x - y = 16 you have to add y to both sides to "move" y to the other side

eviant:

2x+y-y=16+y?

Vocaloid:

good, and simplifying that gives us 2x = 16 + y now subtract 16 from both sides to convert to y = mx + b form then find the slope

eviant:

2x-16=16-16+y

Vocaloid:

good so y = 2x - 16 so the slope of the original line is...?

eviant:

-14+y

Vocaloid:

hm, not quite compare y = mx + b and y = 2x - 16, what is m?

eviant:

2

eviant:

but why did you change it to y=2x-16?

eviant:

shouldn't it have been -14x+y?

Vocaloid:

you cannot combine 2x and 16 to get -14 because they are not like terms

Vocaloid:

it would have to be 2x-16x

Vocaloid:

anyway, now we know the slope is 2, we take the reciprocal which means we write 1/2, then multiply by -1 to get -1/2 as the perpendicular slope = the solution

eviant:

thankx for explaining